# revtex4-2Repair the float

### v3

@@ 4.5 Spectral Functions
Figure 11: The spectral function, $A(\mathbf{k},\omega)$, of the PDW with a half-vortex ($r_{0}=24a_{0}$) evaluated at various energies: $\omega$. From plots (a)-(f) we can see the hole-like character grows with a negative bias and the particle-like portions with positive bias. This redistribution of spectral weight can be used to map out the dispersion.
@@
In this section we analyze the spectral functions of the PDW state in the presence of topological defects, again, paying special attention to the half- vortex. We will also be particularly interested in the plots of the spectral function $A(\mathbf{k},\omega=0)$, defined in equation (4.5), which counts how many quasi-particle states are connected to the ground state within our energy resolution, $\epsilon$. In a metal the spectral function at zero frequency yields the locus of points corresponding to the Fermi surface in the Brillouin zone. In the case of a PDW the spectral function at zero frequency reveals the locus of the Fermi surface of the Bogoliubov quasi-particle. Here we will use the term Bogoliubov Fermi surface to represent the locus of points on the Brillouin zone where there are pockets of Bogoliubov quasi-particle. Since the Bogoliubov quasi-particles are admixtures of electrons and holes different portions of the Bogoliubov Fermi surfaces have electron or hole like character. Still we expect some resemblance between the Fermi surface of the normal state and the Bogoliubov Fermi surfaces of the PDW states; more on this below.
@@
The Bogoliubov Fermi surfaces of the PDW state have been examined in Ref.[25, 46] and in a quasi-1D model of the PDW in Ref.[39]. In these references it was shown that in a time-reversal invariant superconductor, such as the PDW with wave vector $\bm{Q}$, the pockets are separated by gaps in $k$-space, where the SC gap is nonzero, where the condition $\xi_{\mathbf{k}}=\xi_{\mathbf{k}\pm\mathbf{Q}}$ is satisfied; here $\xi_{\mathbf{k}}$ is the quasi-particle dispersion in the non superconducting state. These gaps appear where the Fermi surface of the normal state is perfectly nested and we pair electrons with their time reversed partner.
@@
In this subsection we are interested in the effects of the topological defects on the spectral functions of a PDW. A plot of a few Fermi surfaces of Bogoliubov quasi-particle of the PDW in the presence of a half-vortex can be found in figure 10. Here we consider a half-vortex of $r_{0}=24a_{0}$ possessing a d-wave form factor. Spectral plots of a pristine PDW with both an s-wave and d-wave form factor as well as log plots of the Abrikosov vortex and the double dislocation (both with d-wave form factors) can be found in appendix F figure 16. The latter two plots greatly resemble the spectral function of the pure PDW, small differences only becoming apparent on taking a logarithm. As in the spectral functions of the defect-free PDW of Ref.[25] the portion of the Bogoliubov Fermi surfaces closely resemble “arcs” along the ungapped parts of the normal Fermi surface (i.e. in the absence of the PDW state).
@@
We first observe that the normal state dispersion is partially retained for the presence of the half-vortex. This also holds true for the other defects and is illustrated in appendix F figure 16 where we also overlaid a copy of the normal state Fermi surface with that of the pure PDW (figure 16b). These gapless regions retain the normal state character [48]. Utilizing the weak coupling argument above we realize the modified nesting condition suggests an $s$-wave form factor would also poses these Fermi arcs (see figure 16a).
@@
The most striking feature of the spectral function in a PDW with a half-vortex is the redistribution of spectral weight to regions above the arcs forming discernible loops seen in figure 10a. In figure 10b we zoom in on these loops to indicate a degree of inversion symmetry breaking, seen in the distribution of spectral weight in these loops and along the arcs. Inversion symmetry is broken in the core of the half-vortex where the SC state becomes close to that of an FF state. Note, the formation of these loops does not occur so dramatically for the other two topological defects, but it still happens to some degree. The half-vortex is special in the sense that it couples to both the CDW and the SC degrees-of-freedom, unlike the other two topological defects. The asymmetry in the charge-density induces a significant reshuffling of the spectral weight for the half-vortex according to these plots.
@@
A point worth mentioning at this stage is the apparent coexistence of the “Fermi-Arcs”, just mentioned, and electron/hole pockets [66]. An arc usually refers to a large section of the Fermi surface which is seemingly open-ended. The pockets on the other hand are small closed surfaces. A experimental probe known as Angle-Resolved-Photo-Emission-Spectroscopy (ARPES) can help determine the Fermi surface, but the hole-like regions are invisible to ARPES [67]. We can demonstrate this by looking at log plots of the particle-like Fermi surface (figure 10c) and the hole-like Fermi surface (figure 10d). Here it can be seen the front and backsides of these arcs have primarily particle-like or hole-like character, respectively. Note, this is the case for all the other PDW order parameters as well.
@@
We can further examine the hole-like and particle-like character of certain regions of the Bogoliubov Fermi surface in the presence of the half-vortex by looking at the spectral function for various probing voltages (figures 11a-11f). We see the spectral weight shifts around from one set of loops to another depending on the sign of the bias. Indeed, figure 11 demonstrates a negative bias will grow the hole-like loops of the Fermi surface, while positive biases the electron-like portions. This redistribution of spectral weight pertains to the fact holes are at a negative energy in respect to the Fermi energy and vice versa for the particles.
@@ 5 Discussion and Conclusions
Evidence for the existence of pair density wave superconducting phases (or, at least a PDW component) has continued to grow. In addition to the panoply of evidence in the cuprate superconductors [2] new evidence for PDW order has now been found in other materials such as the heavy fermion superconductor UTe2222 [17, 18], in a monolayer iron superconductor Fe(Te,Se) [19], and in EuRbFe4As4 [20], and in the kagome superconductor CsV3Sb5 [22]. This growing body of evidence of the existence of PDW superconducting states makes the characterization of these phases an important problem.
In this paper we investigated the electronic structure of the BdG Hamiltonian of a unidirectional PDW in two dimensions in the presence of its three topological defects: the half-vortex, the Abrikosov vortex and the double dislocation. In essence we showed that the topological defects of a PDW generically have “halos” which provide evidence for the nature of this superconducting state. However, it is important to distinguish the halos of the PDW topological defects, which occur in the absence of a magnetic field, to the PDW halo of a superconducting vortex which requires the presence of a magnetic field [15, 16, 65]. In contrast, the half-vortex and double dislocation topological defects of the PDW can only be created by impurities.
This work was partly motivated by evidence for half vortices found by Du and collaborators [14] in STM experiments in the superconducting state of the high $T_{c}$ superconductor Bi2Sr2CaCu2O8+δ. The results of our work will also be useful for investigating PDW superconducting state the heavy fermion superconductor $UTe_{2}$. Recent STM experiments in this material have revealed that in its vortices the associated CDW has a dislocation-anti-dislocation dipole structure [68]. Our results provide new ways to investigate the nature of the PDW superconductors by investigating the structure of its interesting topological defects.
@@
The PDW state and its topological defects was treated using the Landau-Ginzburg theory of the Refs. [25, 26, 1] which describes the PDW and its CDW as intertwined orders. In this approach, the CDW order parameter is a composite operator of the two independent PDW order parameter fields. The static configuration of the half-vortex was derived using a non-linear sigma model approximation valid deep in the PDW phase. The BdG Hamiltonian of the PDW was then adapted to include the changes in the PDW order parameter in the presence of the topological defects. The calculation of the electronic states described by the BdG Hamiltonian is not self-consistent. Using the resulting Green functions of our effective theory, obtained numerically, we investigated the $4a_{0}$ CDW of the PDW phase as well as the effects of the topological defects on the electronic states. We should note that this approach is not self-consistent since the PDW order parameter (with or without defects) is fixed. The lack of self-consistency require some caveats on our results that are discussed below.
@@
We analyzed in detail the quasi-particle spectral function of a PDW with a half-vortex defect. In addition to the “arc-like” structure at the Fermi surface of the Bogoliubov quasi-particle states which are seen in the defect-free PDW state [25, 46], we found that the half-vortex induces asymmetric “loop-like” structures above the “arcs”. We attributed the existence of these loops to the breaking of inversion symmetry at the core of the half-vortex. While much of the quasi-particle spectral function is very similar to that of the defect-free PDW, the presence of the half-vortex, and its inversion symmetry breaking, has a clear imprint in the spectral function. We also analyzed the real space position and the voltage dependence of the local differential conductance across a half vortex core. This dependence gives additional evidence for the existence of an FF component of the PDW in the core of the half-vortex.
Since this is not a self-consistent theory, both the spectral functions and the differential tunneling conductance results should be reliable at low energies but cannot be trusted at energies (voltages) substantially higher than the superconducting gap. As a matter of principle we expect that at energies well above the gap the superconducting order parameter should be progressively suppressed and the Bogoliubov quasi-particle effectively should become “normal” electrons. This also implies that the composite order parameters such as the CDW should also be progressively suppressed well above the gap. This is not what happens in our numerics which computed the BdG spectrum in a fixed background of the superconducting order. To have a fully self-consistent theory requires a viable physical mechanism for a PDW which cannot be obtained by a weak coupling BCS-type theory. This is an open problem and a matter of current research.
In most systems in which PDW has been observed it happens in, at best, coexistence with a uniform superconducting state. This happens even in the case of La2-xBaxCuO4 which as, so far, provided the best evidence for PDW order (see Refs.[2, 70] and references therein). Thus it is important to understand what changes are brought about to our results when a PDW coexists with uniform superconducting order parameter. We plan to address this problem in a separate publication.
Contrary to the case of vortices in a superconductor, whose number and separation are controlled by an external magnetic field, impurities are needed to create the half-vortices and double dislocations of a PDW. Here we considered the problem of single isolated topological defects. In practice this will require very clean systems so that the impurities are separated over large distances, larger than the size of the halos. On the other hand, a finite density of disorder has large qualitative effects in states such as the PDW, including the destruction of long-range CDW order [71, 72], and most intriguingly a possible charge-4e superconducting state [44] by proliferation of double dislocations as proposed in Ref.[73]. These important open problems are beyond the scope of this paper.
@@ Acknowledgements.
We thank Ryan Levy, Raman Sohal and Yuxuan Wang for useful discussions pertaining to the setup of the numerical simulation and Vidya Madhavan for addressing issues with the low bias particle-hole symmetry in the STM plots. Furthermore, we thank Peiyan Wu for his help in parallelizing the eigen-solvers, and for correcting the orthogonalization algorithms in these canned routines. MR thanks Mike Stone with his help in clarifying certain aspects of our solution process. This work was supported in part by the US National Science Foundation through the grant NSF DMR 2225920 at the University of Illinois (MR and EF) and by the Sloan Foundation (MR).
@@ Appendix A Bogoliubov-Valatin transformation
We associate each Latin index with position in this section. It is standard to diagonalize this operator using a Bogoliubov-Valatin (BV) transformation, but we will opt for a Nambu formalism instead. In the end we obtain the same BV transformation defining the same quasi-particles; the alternative route of defining our quasi-particle operators before diagonalizing can be found in Ref. [74].
Let’s define the following operator: $\mathbf{\tilde{\psi}}^{T}=[\mathbf{c}_{\uparrow},\mathbf{c}_{\downarrow}]$, where the above components are vectors who consist of electron operators, $\big{[}\mathbf{c}_{\sigma}\big{]}_{i}=\hat{c}_{\mathbf{r}_{i}\sigma}$. Defining our Nambu spinor as $\tilde{\Psi}^{T}=[\tilde{\psi},\tilde{\psi}^{\dagger}]$ we can express our above Hamiltonian as a matrix product
$\hat{H}$ $=\tilde{\Psi}^{\dagger}\mathbf{\tilde{H}}\tilde{\Psi}=\tilde{\Psi}^{\dagger}\Big{(}\mathbf{\tilde{V}}\mathbf{E}\mathbf{\tilde{V}}^{-1}\Big{)}\tilde{\Psi}\equiv\tilde{\gamma}^{\dagger}\mathbf{E}\tilde{\gamma}$ (A.2)
The explicit definitions of the terms presented in equation (A.2) will be covered in the next few paragraphs, but essentially this is just a similarity transformation. The matrix $\mathbf{\tilde{H}}$ takes the following generic form
$\mathbf{\tilde{H}}=\frac{1}{2}\left[\begin{array}[]{c c c c}\mathbf{T}&\mathbf{\tilde{\Delta}}\\ \mathbf{\tilde{\Delta}}^{\dagger}&-\mathbf{T}^{T}\end{array}\right]$ (A.3)
Defining $[\mathbf{t}]_{ij}=-t_{ij}$ and $\mathbf{\Delta}_{\uparrow\uparrow}$/$\mathbf{\Delta}_{\uparrow\downarrow}$ (etc.) for triplet/singlet superconductivity we have the following forms for our sub-matrices
$\mathbf{T}=\left[\begin{array}[]{c c c c}\mathbf{t}&\mathbf{0}\\ \mathbf{0}&\mathbf{t}\end{array}\right],\;\;\mathbf{\tilde{\Delta}}=\left[\begin{array}[]{c c c c}\mathbf{\Delta}_{\uparrow\uparrow}&\mathbf{\Delta}_{\uparrow\downarrow}\\ \mathbf{\Delta}_{\downarrow\uparrow}&\mathbf{\Delta}_{\downarrow\downarrow}\end{array}\right]=-\mathbf{\tilde{\Delta}}^{T}$ (A.4)
@@
We next define the quasi-particle operators, $\tilde{\gamma}=[\mathbf{b}^{\dagger},\mathbf{b}]^{T}$, which can be expressed in terms of the electron creation and annihilation operators with the unitary matrix inducing the similarity transform, $\mathbf{\tilde{V}}$. Our similarity transform takes a very simple form because there are no zero modes in our spectrum due to the finite size of the system (needed for numerical diagonlization). This reveals we have $\pm$ energy pairs, and they are related via complex conjugation of the eigenvalue equation and swapping the top blocks of the BdG equations with the bottom. That is, given (sorted) eigenenergies $E_{l}>0$ ($l\in\{1,\dots,2N^{2}\}$) we have the corresponding negative energy solutions
$\mathbf{\tilde{H}}\begin{bmatrix}\mathbf{\tilde{u}}_{l}\\ \mathbf{\tilde{v}}_{l}\end{bmatrix}=E_{l}\begin{bmatrix}\mathbf{\tilde{u}}_{l}\\ \mathbf{\tilde{v}}_{l}\end{bmatrix}\quad\Rightarrow\quad\mathbf{\tilde{H}}\begin{bmatrix}\mathbf{\tilde{v}}^{*}_{l}\\ \mathbf{\tilde{u}}^{*}_{l}\end{bmatrix}=-E_{l}\begin{bmatrix}\mathbf{\tilde{v}}^{*}_{l}\\ \mathbf{\tilde{u}}^{*}_{l}\end{bmatrix}$ (A.5)
We now define the following $2N^{2}\times 2N^{2}$ matrix $\mathbf{\tilde{u}}$ by setting it’s $l^{th}$ column equal to $\mathbf{\tilde{u}}_{l}$, and similarly for $\mathbf{\tilde{v}}$, so the row index corresponds to the lattice site and the matrix $\big{[}\mathbf{E}\big{]}_{lk}=E_{l}\delta_{lk}$. The unitary matrix inducing the similarity transform takes the following form
$\mathbf{\tilde{V}}=\begin{bmatrix}\mathbf{\tilde{v}}^{*}&\mathbf{\tilde{u}}\\ \mathbf{\tilde{u}}^{*}&\mathbf{\tilde{v}}\end{bmatrix}$ (A.6)
The block matrices, $\mathbf{\tilde{u}}$ and $\mathbf{\tilde{v}}$, contain the real space coherence factors, and provide us with the following electron operators in terms of our quasi-particle operators
$\hat{c}_{i\uparrow}=\tilde{v}^{*}_{il}\hat{b}_{l}^{\dagger}+\tilde{u}_{il}\hat{b}_{l},\;\;\;\;\;\;\hat{c}_{i\uparrow}^{\dagger}=\tilde{v}_{il}\hat{b}_{l}+\tilde{u}^{*}_{il}\hat{b}^{\dagger}_{l}$ (A.7)
$\hat{c}^{\dagger}_{i\downarrow}=\tilde{u}^{*}_{il}\hat{b}_{l}^{\dagger}+\tilde{v}_{il}\hat{b}_{l},\;\;\;\;\;\;\hat{c}_{i\downarrow}=\tilde{u}_{il}\hat{b}_{l}+\tilde{v}^{*}_{il}\hat{b}_{l}^{\dagger}$
@@
|$\ket{G}=\mathcal{N}\exp\Big{(}\frac{1}{2}\hat{c}^{\dagger}_{i\uparrow}S_{ij}\hat{c}^{\dagger}_{j\downarrow}\Big{)}\ket{0}$ (A.10)

### v4

@@ 4.5 Spectral Functions
Figure 11: The spectral function, $A(\mathbf{k},\omega)$, of the PDW with a half-vortex ($r_{0}=24a_{0}$) evaluated at various energies: $\omega$. From plots (a)-(f) we can see the holelike character grows with a negative bias and the particle-like portions with positive bias. This redistribution of spectral weight can be used to map out the dispersion.
@@
In this section we analyze the spectral functions of the PDW state in the presence of topological defects, again, paying special attention to the half- vortex. We will also be particularly interested in the plots of the spectral function $A(\mathbf{k},\omega=0)$, defined in Eq.  (4.5), which counts how many quasiparticle states are connected to the ground state within our energy resolution, $\epsilon$. In a metal the spectral function at zero frequency yields the locus of points corresponding to the Fermi surface in the Brillouin zone. In the case of a PDW the spectral function at zero frequency reveals the locus of the Fermi surface of the Bogoliubov quasiparticle. Here we will use the term Bogoliubov Fermi surface to represent the locus of points on the Brillouin zone where there are pockets of Bogoliubov quasiparticle. Since the Bogoliubov quasiparticles are admixtures of electrons and holes different portions of the Bogoliubov Fermi surfaces have electron or holelike character. Still we expect some resemblance between the Fermi surface of the normal state and the Bogoliubov Fermi surfaces of the PDW states; we give more on this below.
@@
The Bogoliubov Fermi surfaces of the PDW state have been examined in Refs. [25, 46] and in a quasi-1D model of the PDW in Ref. [39]. In these references it was shown that in a time-reversal invariant superconductor, such as the PDW with wave vector $\bm{Q}$, the pockets are separated by gaps in $k$-space, where the SC gap is nonzero, where the condition $\xi_{\mathbf{k}}=\xi_{\mathbf{k}\pm\mathbf{Q}}$ is satisfied; here $\xi_{\mathbf{k}}$ is the quasiparticle dispersion in the non superconducting state. These gaps appear where the Fermi surface of the normal state is perfectly nested and we pair electrons with their time reversed partner.
@@
In this subsection we are interested in the effects of the topological defects on the spectral functions of a PDW. A plot of a few Fermi surfaces of Bogoliubov quasiparticle of the PDW in the presence of a half-vortex can be found in Fig. 10. Here we consider a half-vortex of $r_{0}=24a_{0}$ possessing a $d$-wave form factor. Spectral plots of a pristine PDW with both an $s$-wave and $d$-wave form factor as well as log plots of the Abrikosov vortex and the double dislocation (both with $d$-wave form factors) can be found in Appendix F, Fig. 16. The latter two plots greatly resemble the spectral function of the pure PDW, small differences only becoming apparent on taking a logarithm. As in the spectral functions of the defect-free PDW of Ref.[25] the portion of the Bogoliubov Fermi surfaces closely resemble “arcs” along the ungapped parts of the normal Fermi surface (i.e., in the absence of the PDW state).
@@
We first observe that the normal state dispersion is partially retained for the presence of the half-vortex. This also holds true for the other defects and is illustrated in Appendix F, Fig. 16, where we also overlaid a copy of the normal state Fermi surface with that of the pure PDW [Fig. 16b]. These gapless regions retain the normal state character [48]. Utilizing the weak coupling argument above we realize the modified nesting condition suggests an $s$-wave form factor would also possess these Fermi arcs (see Fig. 16a).
@@
The most striking feature of the spectral function in a PDW with a half-vortex is the redistribution of spectral weight to regions above the arcs forming discernible loops seen in Fig. 10a. In Fig. 10b we zoom in on these loops to indicate a degree of inversion symmetry breaking, seen in the distribution of spectral weight in these loops and along the arcs. Inversion symmetry is broken in the core of the half-vortex where the SC state becomes close to that of an FF state. Note that the formation of these loops does not occur so dramatically for the other two topological defects, but it still happens to some degree. The half-vortex is special in the sense that it couples to both the CDW and the SC degrees-of-freedom, unlike the other two topological defects. The asymmetry in the charge density induces a significant reshuffling of the spectral weight for the half-vortex according to these plots.
@@
A point worth mentioning at this stage is the apparent coexistence of the “Fermi-Arcs”, just mentioned, and electron/hole pockets [65]. An arc usually refers to a large section of the Fermi surface which is seemingly open-ended. The pockets on the other hand are small closed surfaces. An experimental probe known as angle-resolved photo emission spectroscopy (ARPES) can help determine the Fermi surface, but the holelike regions are invisible to ARPES [66]. We can demonstrate this by looking at log plots of the particle-like Fermi surface [Fig. 10c] and the holelike Fermi surface [Fig. 10d]. Here it can be seen that the front and backsides of these arcs have primarily particle-like or holelike character, respectively. Note that this is the case for all the other PDW order parameters as well.
@@
We can further examine the holelike and particlelike character of certain regions of the Bogoliubov Fermi surface in the presence of the half-vortex by looking at the spectral function for various probing voltages [Figs. 11a-11f]. We see the spectral weight shifts around from one set of loops to another depending on the sign of the bias. Indeed, Fig. 11 demonstrates a negative bias will grow the holelike loops of the Fermi surface, while positive biases the electron-like portions. This redistribution of spectral weight pertains to the fact holes are at a negative energy in respect to the Fermi energy and vice versa for the particles.
@@ 5 Discussion and Conclusions
Evidence for the existence of pair density wave superconducting phases (or, at least a PDW component) has continued to grow. In addition to the panoply of evidence in the cuprate superconductors [2] new evidence for PDW order has now been found in other materials such as the heavy fermion superconductor $UTe_{2}$ [17, 18], in a monolayer iron superconductor Fe(Te,Se) [19], in EuRbFe4As4 [20], and in the kagome superconductor CsV3Sb5 [22]. This growing body of evidence of the existence of PDW superconducting states makes the characterization of these phases an important problem.
In this paper we investigated the electronic structure of the BdG Hamiltonian of a unidirectional PDW in two dimensions in the presence of its three topological defects: the half-vortex, the Abrikosov vortex, and the double dislocation. In essence we showed that the topological defects of a PDW generically have “halos” which provide evidence for the nature of this superconducting state. However, it is important to distinguish the halos of the PDW topological defects, which occur in the absence of a magnetic field, to the PDW halo of a superconducting vortex which requires the presence of a magnetic field [15, 16, 64]. In contrast, the half-vortex and double dislocation topological defects of the PDW can be created only by impurities.
This work was partly motivated by evidence for half vortices found by Du and collaborators [14] in STM experiments in the superconducting state of the high $T_{c}$ superconductor Bi2Sr2CaCu2O8+δ. The results of our work will also be useful for investigating in the PDW superconducting state the heavy fermion superconductor $UTe_{2}$. Recent STM experiments in this material have revealed that in its vortices the associated CDW has a dislocation-anti-dislocation dipole structure [67]. Our results provide alternative ways to investigate the nature of the PDW superconductors by investigating the structure of its interesting topological defects.
@@
The PDW state and its topological defects was treated using the Landau-Ginzburg theory of the Refs. [25, 26, 1], which describes the PDW and its CDW as intertwined orders. In this approach, the CDW order parameter is a composite operator of the two independent PDW order parameter fields. The static configuration of the half-vortex was derived using a non-linear sigma model approximation valid deep in the PDW phase. The BdG Hamiltonian of the PDW was then adapted to include the changes in the PDW order parameter in the presence of the topological defects. The calculation of the electronic states described by the BdG Hamiltonian is not self-consistent. Using the resulting Green functions of our effective theory, obtained numerically, we investigated the $4a_{0}$ CDW of the PDW phase as well as the effects of the topological defects on the electronic states. We should note that this approach is not self-consistent since the PDW order parameter (with or without defects) is fixed. The lack of self-consistency require some caveats on our results that are discussed below.
@@
We analyzed in detail the quasiparticle spectral function of a PDW with a half-vortex defect. In addition to the “arc-like” structure at the Fermi surface of the Bogoliubov quasiparticle states which are seen in the defect-free PDW state [25, 46], we found that the half-vortex induces asymmetric “loop-like” structures above the “arcs”. We attributed the existence of these loops to the breaking of inversion symmetry at the core of the half-vortex. While much of the quasiparticle spectral function is very similar to that of the defect-free PDW, the presence of the half-vortex, and its inversion symmetry breaking, has a clear imprint in the spectral function. We also analyzed the real space position and the voltage dependence of the local differential conductance across a half vortex core. This dependence gives additional evidence for the existence of an FF component of the PDW in the core of the half-vortex.
Since this is not a self-consistent theory, both the spectral functions and the differential tunneling conductance results should be reliable at low energies but cannot be trusted at energies (voltages) substantially higher than the superconducting gap. As a matter of principle we expect that at energies well above the gap the superconducting order parameter should be progressively suppressed and the Bogoliubov quasiparticle effectively should become “normal” electrons. This also implies that the composite order parameters such as the CDW should also be progressively suppressed well above the gap. This is not what happens in our numerics which computed the BdG spectrum in a fixed background of the superconducting order. To have a fully self-consistent theory requires a viable physical mechanism for a PDW which cannot be obtained by a weak coupling BCS-type theory. This is an open problem and a matter of current research.
In most systems in which PDW has been observed it happens in, at best, coexistence with a uniform superconducting state. This happens even in the case of La2-xBaxCuO4, which has, so far, provided the best evidence for PDW order (see Refs. [2, 69] and references therein). Thus it is important to understand what changes are brought about to our results when a PDW coexists with uniform superconducting order parameter. We plan to address this problem in a separate publication.
[…]
@@ Acknowledgements.
We thank Ryan Levy, Raman Sohal, and Yuxuan Wang for useful discussions pertaining to the setup of the numerical simulation and Vidya Madhavan for addressing issues with the low bias particle-hole symmetry in the STM plots. Furthermore, we thank Peiyan Wu for his help in parallelizing the eigen-solvers, and for correcting the orthogonalization algorithms in these canned routines. M.R. thanks Mike Stone with his help in clarifying certain aspects of our solution process. This work was supported in part by the US National Science Foundation through Grant No. NSF DMR 2225920 at the University of Illinois (M.R. and E.F.) and by the (2017) Sloan Foundation (M.R.).
@@ Appendix A Bogoliubov-Valatin transformation
We associate each Latin index with position in this appendix. It is standard to diagonalize this operator using a Bogoliubov-Valatin (BV) transformation, but we will opt for a Nambu formalism instead. In the end we obtain the same BV transformation defining the same quasiparticles; the alternative route of defining our quasiparticle operators before diagonalizing can be found in Ref. [73].
Let’s define the following operator: $\mathbf{\tilde{\psi}}^{T}=[\mathbf{c}_{\uparrow},\mathbf{c}_{\downarrow}]$, where the above components are vectors that consist of electron operators, $\big{[}\mathbf{c}_{\sigma}\big{]}_{i}=\hat{c}_{\mathbf{r}_{i}\sigma}$. Defining our Nambu spinor as $\tilde{\Psi}^{T}=[\tilde{\psi},\tilde{\psi}^{\dagger}]$ we can express our above Hamiltonian as a matrix product:
$\hat{H}$ $=\tilde{\Psi}^{\dagger}\mathbf{\tilde{H}}\tilde{\Psi}=\tilde{\Psi}^{\dagger}\Big{(}\mathbf{\tilde{V}}\mathbf{E}\mathbf{\tilde{V}}^{-1}\Big{)}\tilde{\Psi}\equiv\tilde{\gamma}^{\dagger}\mathbf{E}\tilde{\gamma}.$ (A.2)
The explicit definitions of the terms presented in Eq.  (A.2) will be covered in the next few paragraphs, but essentially this is just a similarity transformation. The matrix $\mathbf{\tilde{H}}$ takes the following generic form:
$\mathbf{\tilde{H}}=\frac{1}{2}\left[\begin{array}[]{c c c c}\mathbf{T}&\mathbf{\tilde{\Delta}}\\ \mathbf{\tilde{\Delta}}^{\dagger}&-\mathbf{T}^{T}\end{array}\right].$ (A.3)
Defining $[\mathbf{t}]_{ij}=-t_{ij}$ and $\mathbf{\Delta}_{\uparrow\uparrow}$/$\mathbf{\Delta}_{\uparrow\downarrow}$ (etc.) for triplet/singlet superconductivity we have the following forms for our sub matrices:
$\mathbf{T}=\left[\begin{array}[]{c c c c}\mathbf{t}&\mathbf{0}\\ \mathbf{0}&\mathbf{t}\end{array}\right],\;\;\mathbf{\tilde{\Delta}}=\left[\begin{array}[]{c c c c}\mathbf{\Delta}_{\uparrow\uparrow}&\mathbf{\Delta}_{\uparrow\downarrow}\\ \mathbf{\Delta}_{\downarrow\uparrow}&\mathbf{\Delta}_{\downarrow\downarrow}\end{array}\right]=-\mathbf{\tilde{\Delta}}^{T}.$ (A.4)
@@
We next define the quasiparticle operators, $\tilde{\gamma}=[\mathbf{b}^{\dagger},\mathbf{b}]^{T}$, which can be expressed in terms of the electron creation and annihilation operators with the unitary matrix inducing the similarity transform, $\mathbf{\tilde{V}}$. Our similarity transform takes a very simple form because there are no zero modes in our spectrum due to the finite size of the system (needed for numerical diagonlization). This reveals we have $\pm$ energy pairs, and they are related via complex conjugation of the eigenvalue equation and swapping the top blocks of the BdG equations with the bottom. That is, given (sorted) eigenenergies $E_{l}>0$ ($l\in\{1,\dots,2N^{2}\}$) we have the corresponding negative energy solutions
$\mathbf{\tilde{H}}\begin{bmatrix}\mathbf{\tilde{u}}_{l}\\ \mathbf{\tilde{v}}_{l}\end{bmatrix}=E_{l}\begin{bmatrix}\mathbf{\tilde{u}}_{l}\\ \mathbf{\tilde{v}}_{l}\end{bmatrix}\quad\Rightarrow\quad\mathbf{\tilde{H}}\begin{bmatrix}\mathbf{\tilde{v}}^{*}_{l}\\ \mathbf{\tilde{u}}^{*}_{l}\end{bmatrix}=-E_{l}\begin{bmatrix}\mathbf{\tilde{v}}^{*}_{l}\\ \mathbf{\tilde{u}}^{*}_{l}\end{bmatrix}.$ (A.5)
We now define the following $2N^{2}\times 2N^{2}$ matrix $\mathbf{\tilde{u}}$ by setting it’s $l^{th}$ column equal to $\mathbf{\tilde{u}}_{l}$, and similarly for $\mathbf{\tilde{v}}$, so the row index corresponds to the lattice site and the matrix $\big{[}\mathbf{E}\big{]}_{lk}=E_{l}\delta_{lk}$. The unitary matrix inducing the similarity transform takes the following form:
$\mathbf{\tilde{V}}=\begin{bmatrix}\mathbf{\tilde{v}}^{*}&\mathbf{\tilde{u}}\\ \mathbf{\tilde{u}}^{*}&\mathbf{\tilde{v}}\end{bmatrix}.$ (A.6)
The block matrices, $\mathbf{\tilde{u}}$ and $\mathbf{\tilde{v}}$, contain the real space coherence factors, and provide us with the following electron operators in terms of our quasiparticle operators:
$\hat{c}_{i\uparrow}=\tilde{v}^{*}_{il}\hat{b}_{l}^{\dagger}+\tilde{u}_{il}\hat{b}_{l},\;\;\;\;\;\;\hat{c}_{i\uparrow}^{\dagger}=\tilde{v}_{il}\hat{b}_{l}+\tilde{u}^{*}_{il}\hat{b}^{\dagger}_{l},$ (A.7)
$\hat{c}^{\dagger}_{i\downarrow}=\tilde{u}^{*}_{il}\hat{b}_{l}^{\dagger}+\tilde{v}_{il}\hat{b}_{l},\;\;\;\;\;\;\hat{c}_{i\downarrow}=\tilde{u}_{il}\hat{b}_{l}+\tilde{v}^{*}_{il}\hat{b}_{l}^{\dagger}.$
@@
[…]